OpenStudy (anonymous):

[4.07]What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (lgbasallote):

it says it's parallel to the line that passes through (3,6) and (1,-2) so just get the slope of that. \[\implies \frac{y_2 - y_1}{x_2 - x_1} \implies \frac{-2 - 6}{1 - 3} \implies \frac{-8}{-2} \implies 4\] that's it. as simple as that. got it?

OpenStudy (anonymous):

thankss

OpenStudy (lgbasallote):

welcome